OpenStudy (anonymous):

Which property justifies that 4*37*25=4*25*37

jimthompson5910 (jim_thompson5910):

Hint: which property states that x * y = y * x ??

OpenStudy (anonymous):

umm identity

jimthompson5910 (jim_thompson5910):

no the identity properties are x + 0 = x 0 + x = x x * 1 = x 1 * x = x

jimthompson5910 (jim_thompson5910):

basically, doing some operation to x, doesn't change x

OpenStudy (anonymous):

oh is it associative

jimthompson5910 (jim_thompson5910):

no that says that if you have 3 numbers x, y and z, then x + (y+z) = (x+y)+z x * (y*z) = (x*y)*z

jimthompson5910 (jim_thompson5910):

you're re-associating and moving around parenthesis

OpenStudy (anonymous):

communtive

jimthompson5910 (jim_thompson5910):

commutative property of ____

OpenStudy (anonymous):

so it has to be inverse then

jimthompson5910 (jim_thompson5910):

commutative property of inverse no such property

jimthompson5910 (jim_thompson5910):

is it commutative property of addition or commutative property of multiplication

OpenStudy (anonymous):

haha of i thought communtive was wrong

jimthompson5910 (jim_thompson5910):

oh, no commutative property is (partially) correct

jimthompson5910 (jim_thompson5910):

just not complete since there are 2 properties that are commutative

OpenStudy (anonymous):

commutative property of multiplication

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

yw